# Bayesian Despeckling of Structured Sources

### v1

@@ 2 Bayesian Despeckling via QMAP
The cost function in (3) comprises of two terms. The first term represents a fidelity criterion derived from the negative likelihood of the observations given the signal under the multiplicative noise model. Note that given $X^{n}=x^{n}$, $Y^{n}\sim\mathcal{N}({\bf 0},X^{2})$, where $X=\mathop{\rm Diag}\nolimits({\bf x})$. Therefore,
@@
A key advantage of this approach is that it leads to a tractable approach for modeling and utilizing the sources structure. It can be shown that for structured sources, i.e., the sources with information dimension strictly less than one [28, 29], identified by having singularities in their distributions, e.g., spike and slap distribution, the weights ${\bf w}$ can be divided into two groups: 1) A small set of weights that identify the key structure of the source, and 2) a large set of weights that have negligible impact in the optimization [27].
@@
Solving the BD-QMAP optimization is challenging due to the nature of the cost function, which combines a non-convex loss term derived from the log-likelihood and a regularization term defined on the discretized space of sequences. To address this complexity, the optimization can be simplified by restricting the search space to $\mathcal{X}_{b}^{n}$, the space of quantized sequences. As explained in Section 5.1, this restriction enables the application of the Viterbi algorithm for efficient optimization. While this approach is suboptimal compared to solving the original problem, it offers significant computational advantages. Moreover, for cases such as the piecewise-constant source studied later, this simplification facilitates learning the structure of the jumps, allowing the original optimization to be solved in the continuous space with improved computational efficiency.
@@ 3.1 Structured memoryless source
Consider an i.i.d process ${\bf X}$ such that $X_{i}\sim(1-q_{0})\delta_{x_{m}}+q_{0}{\rm Uniform}(x_{m},x_{M})$, Setting $k=1$ and using the results of [27, Sec. 3.1] to simplfy the second term in (3), the BD-QMAP algorithm can be written as
@@ 3.2 Piecewise constant source
where for any $k\in\{0,\ldots,n-1\}$, the inner optimization is over $(n_{1},\ldots,n_{k+1})\in(\mathds{N}^{+})^{k+1}$, such that $\sum_{i=1}^{k+1}n_{i}=n$. Moreover, given ${\bf n}=(n_{1},\ldots,n_{k+1})$, $\sum_{j=1}^{k+1}n_{j}=n$, for any $j\in\{1,\dots,k+1\}$,
$\mathcal{I}_{j}({\bf n})=\{\sum_{i=1}^{j-1}n_{i}+1,\ldots,\sum_{i=1}^{j}n_{i}\}.$ (7)
Moreover, for $i\in\mathcal{I}_{j}$,
@@
In other words, Lemma 1 implies that solving the BD-QMAP algorithm is equivalent to identifying the locations of the jumps such that, over each constant interval, the source input is estimated through appropriate averaging of the noisy observations, thereby minimizing the corresponding loss function. It is important to emphasize that if the regularization term enforcing the source structure were absent (i.e., if $\lambda=0$), the concavity of the $\log$ function would result in the solution equaling the absolute value of input noisy sequence, with the maximal number of jumps. On the other hand, choosing $\lambda$ too larges, will ensure that the output has no jumps.
@@
Consider $X^{n}$ generated by a stationary first-order Markov source $\mathbf{X}=\{X_{i}\}_{i\geq 1}$, characterized by $p(X_{i+1}=x_{i+1}|X_{i}=x_{i})=(1-q_{0})\delta_{x_{i}}+q_{0}{\rm\pi_{c}}(x_{m},x_{M})$, where $\pi_{c}$ denotes the pdf of an absolutely continuous distribution with bounded support $[x_{m},x_{M}]$, $x_{m}>0$. Let $\operatorname{E}[X_{i}^{2}]=\eta^{2}$. Let ${\hat{X}}^{n}$ denote the solution of BD-QMAP optimization, when the number of jumps are known to be $k=k(n)$. Then,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
$-2M^{2}\Big{(}{\epsilon\over 1-\epsilon}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)},$
where $T\sim\operatorname{Geometric}(q_{0})$ is representing the constant interval lengthes in the source.
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}(3-\sqrt{2/\pi}-\sqrt{2\pi})-q_{0}^{2}(2-\sqrt{2\pi})+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])+\upsilon_{n},$
@@ 4 Proofs
Let $T_{i}\overset{iid}{\sim}\operatorname{Geometric}(q_{0})\quad\forall i\in\{1,\dots,N\}$ , then for any $t>0$
$P\left(\left|\frac{\sum_{i=1}^{N}T_{i}}{N}-\frac{1}{q_{0}}\right|>t\right)\leq 2e^{-N\left(q_{0}t+\ln\frac{1}{1+q_{0}t}\right)}.$
Proof of Lemma 1.
Note that the set $\mathcal{U}^{n}$ can be partitioned into $n$ subsets depending on the number of jumps. That is, $\mathcal{U}^{n}=\bigcup_{i=0}^{n-1}\mathcal{U}_{k},$ where, for $k=0,\ldots,n-1$,
$\mathcal{U}_{k}=\{u^{n}\in\mathcal{U}^{n}:\;\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}=k\}.$
@@
$\min_{k\in\{0,\ldots,n-1\}}\min_{u^{n}\in\mathcal{U}_{k}}[{1\over n}\sum_{j=1}^{n}(\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}})+{1\over n-1}(\lambda+{1\over b}\eta)\sum_{i=2}^{n}\mathbbmss{1}_{u_{i}\neq u_{i-1}}]$
@@
which yields the desired result. ∎
Proof of Theorem 1.
@@
$\stackrel{{\scriptstyle\rm(a)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{k\over n}-q_{0}|{M^{2}\over k}\sum_{j=1}^{k}T_{j}(1+{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2})$
$\stackrel{{\scriptstyle\rm(b)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{1\over q_{0}}-{n\over k}|M^{2}q_{0}(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2}),$ (15)
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]$ $\geq q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2})\right]$
$=q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right],$ (16)
where the last line follows because the source $\mathbb{X}$ and the speckle noise process are independent. We first focus on the first term. Note that
@@
$2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right]$ $\geq 2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}]$ (21)
@@
$+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])-2M^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right].$ (23)
@@
$\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right]$ $=\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(\mathbbmss{1}_{\mathcal{E}}+\mathbbmss{1}_{\mathcal{E}^{c}})\right]$
@@
$P(k(n)\geq nq_{0}(1+\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}<n)$
$=P\left({1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}-{1\over q_{0}}<{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}-{1\over q_{0}}\right)$
$\leq P\left(|{1\over q_{0}}-{1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}|>{1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}\right)$
@@
$P(k(n)\leq nq_{0}(1-\epsilon))$ $=P(\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}>n)$
$=P({1\over\lfloor nq_{0}(1-\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}-{1\over q_{0}}>{n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}})$
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$ (28)
$-2M^{2}q_{0}\Big{(}{\epsilon\over q_{o}(1-\epsilon)}+n{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+n{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)}.$ (29)
Proof of Corollary 1.

### v2

@@ 2 Bayesian Despeckling via QMAP
The cost function in (3) comprises of two terms. The first term represents a fidelity criterion derived from the negative likelihood of the observations given the signal under the multiplicative noise model. Note that given $X^{n}=x^{n}$, $Y^{n}\sim\mathcal{N}({\bf 0},X^{2})$, where $X=\mathop{\rm Diag}\nolimits(x^{n})$. Therefore,
@@
A key advantage of this approach is that it leads to a tractable approach for modeling and utilizing the source’s structure. It can be shown that for structured sources, i.e., the sources with information dimension strictly less than one [28, 29], identified by having singularities in their distributions, e.g., spike and slap distribution, the weights ${\bf w}$ can be divided into two groups: 1) A small set of weights that identify the key structure of the source, and 2) a large set of weights that have negligible impact in the optimization [27].
@@
Solving the BD-QMAP optimization is challenging due to the nature of the cost function, which combines a non-convex loss term derived from the log-likelihood and a regularization term defined on the discretized space of sequences. To address this complexity, the optimization can be simplified by restricting the search space to $\mathcal{X}_{b}^{n}$, the space of quantized sequences. As will be explained in Section 5.1, this restriction enables the application of the Viterbi algorithm for efficient optimization. While this approach is suboptimal compared to solving the original problem, it offers significant computational advantages. Moreover, for cases such as the piecewise-constant source studied later, this simplification facilitates learning the structure of the jumps, allowing the original optimization to be solved in the continuous space with improved computational efficiency.
@@ 3.1 Structured memoryless source
Consider an i.i.d. process ${\bf X}$ such that $X_{i}\sim(1-q_{0})\delta_{x_{m}}+q_{0}{\rm Uniform}(x_{m},x_{M})$, Setting $k=1$ and using the results of [27, Sec. 3.1] to simplify the second term in (3), the BD-QMAP algorithm can be written as
@@ 3.2 Piecewise constant source
Consider ${\bf n}=(n_{1},\ldots,n_{k+1})\in(\mathbb{N}^{+})^{k+1}$, such that $\sum_{j=1}^{k+1}n_{j}=n$. Then, for any $j\in\{1,\dots,k+1\}$, define
$\mathcal{I}_{j}({\bf n})=\Big{\{}\sum_{i=1}^{j-1}n_{i}+1,\ldots,\sum_{i=1}^{j}n_{i}\Big{\}}.$ (6)
@@
where for any $k\in\{0,\ldots,n-1\}$, the inner optimization is over $(n_{1},\ldots,n_{k+1})\in(\mathbb{N}^{+})^{k+1}$, such that $\sum_{i=1}^{k+1}n_{i}=n$. Moreover, for $i\in\mathcal{I}_{j}({\bf n})$,
@@
In other words, Lemma 1 implies that solving the BD-QMAP algorithm is equivalent to identifying the locations of the jumps such that, over each constant interval, the source input is estimated through appropriate averaging of the noisy observations, thereby minimizing the corresponding loss function. It is important to emphasize that if the regularization term enforcing the source structure were absent (i.e., if $\lambda=0$), the concavity of the $\log$ function would result in the solution equaling the absolute value of input noisy sequence, with the maximal number of jumps. On the other hand, choosing $\lambda$ too large, will ensure that the output has no jumps.
@@
Consider $X^{n}$ generated by a stationary first-order Markov source $\mathbf{X}=\{X_{i}\}_{i\geq 1}$, characterized by $p(X_{i+1}=x_{i+1}|X_{i}=x_{i})=(1-q_{0})\delta_{x_{i}}+q_{0}{\rm\pi_{c}}(x_{m},x_{M})$, where $\pi_{c}$ denotes the pdf of an absolutely continuous distribution with bounded support $[x_{m},x_{M}]$, $x_{m}>0$. Let ${\hat{X}}^{n}$ denote the solution of BD-QMAP optimization, when the number of jumps are known to be $k=k(n)$. Then,
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq 2q_{0}\eta^{2}\Big{(}q_{0}c_{1}+q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]\Big{)}-\upsilon_{n},$ (9)
where $\eta^{2}=\operatorname{E}[X_{i}^{2}]$, $c_{1}=3-\sqrt{2/\pi}-2\sqrt{\pi}$ and $c_{2}=\sqrt{\pi}-2$, and $T\sim\operatorname{Geometric}(q_{0})$. Here,
$\upsilon_{n}=2x_{M}^{2}\Big{(}{\epsilon\over 1-\epsilon}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+nq_{0}{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)},$
with $t_{1}={1/q_{0}}-{n/\lfloor nq_{0}(1+\epsilon)\rfloor}$ and $t_{2}={n/\lfloor nq_{0}(1-\epsilon)\rfloor}-{1/q_{0}}$.
The following corollary, presents a simplified form of Theorem 1, for large values of $n$.
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}(q_{0}c_{1}-q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])-\upsilon_{n},$
@@
Finally, to better understand the term in (9) that involves expectation over $T$, note that for small values of $q_{0}$, with high probability, $T$ is large. On the other hand, for large values of $T$, $f(T)=1/4+O(T^{-1})$, as
$f(T)=T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})$ $=T(1-\exp(-\frac{1}{2}+\frac{T}{2}\ln(1-\frac{1}{T})-\frac{T-1}{2}\ln(1-\frac{2}{T})))$
$=T(1-\exp(-\frac{1}{2}+\frac{T}{2}(-\frac{1}{T}-\frac{1}{2T^{2}}+O(T^{-3}))-\frac{T-1}{2}(-\frac{2}{T}-\frac{2}{T^{2}}+O(T^{-3}))))$
$=T(1-\exp(-\frac{1}{4T}-\frac{1}{T^{2}}+O(T^{-2})))$
$=T(1-1+\frac{1}{4T}+O(T^{-2}))$
$=\frac{1}{4}+O(T^{-1}).$
@@ 4 Proofs
Assume that $T_{i}\overset{iid}{\sim}\operatorname{Geometric}(q_{0})$, $i=1,\ldots,N$. Then for any $t>0$
$\operatorname{P}\left(\left|\frac{\sum_{i=1}^{N}T_{i}}{N}-\frac{1}{q_{0}}\right|>t\right)\leq 2e^{-N\left(q_{0}t+\ln\frac{1}{1+q_{0}t}\right)}.$
4.1 Proof of Lemma 1
Note that the set $\mathcal{U}^{n}$ can be partitioned into $n$ subsets depending on the number of jumps. That is, $\mathcal{U}^{n}=\bigcup_{k=0}^{n-1}\mathcal{U}_{k},$ where, for $k=0,\ldots,n-1$,
$\mathcal{U}_{k}=\{u^{n}\in\mathcal{U}^{n}:\;\sum_{i=2}^{n}\mathds{1}_{u_{i}\neq u_{i-1}}=k\}.$
@@ 4.1 Proof of Lemma 1
$\min_{k\in\{0,\ldots,n-1\}}\min_{u^{n}\in\mathcal{U}_{k}}[{1\over n}\sum_{j=1}^{n}(\log u_{i}^{2}+{Y_{i}^{2}\over u_{i}^{2}})+{1\over n-1}(\lambda+{1\over b}\eta)\sum_{i=2}^{n}\mathds{1}_{u_{i}\neq u_{i-1}}]$
@@
which yields the desired result.
4.2 Proof of Theorem 1
@@
$\stackrel{{\scriptstyle\rm(a)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{k\over n}-q_{0}|{x_{M}^{2}\over k}\sum_{j=1}^{k}T_{j}(1+{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2})$
$\stackrel{{\scriptstyle\rm(b)}}{{\geq}}{q_{0}\over k}\sum_{j=1}^{k}\alpha_{j}^{2}T_{j}\left(1-\sqrt{{1\over T_{j}}\sum_{l\in\mathcal{I}_{j}}W_{l}^{2}}\right)^{2}-|{1\over q_{0}}-{n\over k}|x_{M}^{2}q_{0}(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2}),$ (16)
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n,{\rm(ML)}}\|_{2}^{2}]$ $\geq q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(1+{1\over n}\sum_{l=1}^{n}W_{l}^{2})\right]$
$=q_{0}\eta^{2}\operatorname{E}\left[T\Big{(}1-\sqrt{{1\over T}\sum_{i=1}^{T}W_{i}^{2}}\;\Big{)}^{2}\right]-2x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right],$ (17)
where the last line follows because the source $\mathbf{X}$ and the speckle noise process are independent. We first focus on the first term. Note that
@@
$2\operatorname{E}\left[T-\sqrt{2T}{\Gamma({T+1\over 2})\over\Gamma({T\over 2})}\right]$ $\geq 2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]$ (22)
@@
$+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])-2x_{M}^{2}q_{0}\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right].$ (24)
@@
$\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|\right]$ $=\operatorname{E}\left[|{1\over q_{0}}-{n\over k}|(\mathds{1}_{\mathcal{E}}+\mathds{1}_{\mathcal{E}^{c}})\right]$
@@
$\operatorname{P}(k(n)\geq nq_{0}(1+\epsilon))$ $=\operatorname{P}(\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}<n)$
$=\operatorname{P}\left({1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}-{1\over q_{0}}<{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}-{1\over q_{0}}\right)$
$\leq\operatorname{P}\left(|{1\over q_{0}}-{1\over\lfloor nq_{0}(1+\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1+\epsilon)\rfloor}T_{i}|>{1\over q_{0}}-{n\over\lfloor nq_{0}(1+\epsilon)\rfloor}\right)$
@@
$\operatorname{P}(k(n)\leq nq_{0}(1-\epsilon))$ $=\operatorname{P}(\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}>n)$
$=\operatorname{P}({1\over\lfloor nq_{0}(1-\epsilon)\rfloor}\sum_{i=1}^{\lfloor nq_{0}(1-\epsilon)\rfloor}T_{i}-{1\over q_{0}}>{n\over\lfloor nq_{0}(1-\epsilon)\rfloor}-{1\over q_{0}})$
@@
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}])$ (29)
$-2x_{M}^{2}q_{0}\Big{(}{\epsilon\over q_{o}(1-\epsilon)}+n{\rm e}^{-\lfloor nq_{0}(1+\epsilon)\rfloor\left(q_{0}t_{1}+\ln\frac{1}{1+q_{0}t_{1}}\right)}+n{\rm e}^{-\lfloor nq_{0}(1-\epsilon)\rfloor\left(q_{0}t_{2}+\ln\frac{1}{1+q_{0}t_{2}}\right)}\Big{)}.$ (30)
4.3 Proof of Corollary 1